# Quantum Optimization Benchmark Library The Intractable Decathlon

### v1

@@ 6 Conclusion and Discussion
It is widely believed that quantum computers can achieve an advantage over classical computers for some optimization problem instances. While complexity theory suggests that, for NP-hard problems, the speed-up for finding provably optimal solutions is in general at most a quadratic [13, 65, 58, 97], better quantum approximation algorithms with an exponential speed-up [123] or quantum optimization heuristics that outperform classical algorithms on certain instances may still exist. However, practical demonstrations of quantum advantage in optimization are still missing.
To identify problems where quantum advantage might be possible and to track the progress to get there, systematic analysis and benchmarking is key. To this end, we introduce QOBLIB, an open-source quantum optimization benchmarking library [193]. We present ten classes of combinatorial optimization problems with instances of varying size and complexity for benchmarking classical and quantum optimization algorithms. Further, we define a reporting standard that ensures reproducibility and fair comparison while remaining flexible for evaluating different algorithms and hardware platforms.
Most problems are challenging for state-of-the-art classical solvers even at moderate size and many of them are related to use cases of practical relevance. The problem classes exhibit diverse properties, such as different densities and types of constraints, to enable the study of different mathematical formulations and algorithmic approaches. This is essential for identifying problem structures suitable for quantum advantage in optimization.
We provide classical baseline results for all problem classes and quantum baseline results for selected ones to kick off this effort. However, performing exhaustive benchmarks is a massive endeavor and can only be achieved by the broader scientific community. Thus, we invite all researchers interested in advancing (quantum) optimization to investigate the provided problem instances and submit solutions using algorithms and hardware of their choice. We also want to emphasize the importance of improving classical methodologies, amongst other reasons, because quantum advantage can only be claimed if the best-known classical methods are outperformed. To accelerate the overall research, we also encourage researchers to report results for approaches that did not work well. To conclude, QOBLIB is intended to be a living repository: novel solutions will be included on a rolling basis, and novel problem instances or adaptions of existing ones may be added as the state-of-the-art evolves.
@@ Acknowledgments
We want to thank Karen Aardal at TU-Delft for providing results for the Market Split instances using lattice point enumeration and Nils-Christian Kempke at ZIB for providing results by GPU-accellerated Schroeppel Shamir enumeration. We would like to thank Sébastien Designolle at the Zuse Institute Berlin for useful discussions on the Blended FW algorithm used in the classical benchmarking of the minimum Birkhoff decomposition problem (Section 4.3.4). We also thank Julien Gacon for valuable discussions and a helpful review and Takashi Imamichi for assistance with software issues, sharing his insights, and providing feedback. Additionally, we are grateful to Andrew Wack and Paul Nation for insightful conversations about the benchmarking of quantum algorithms and their feedback on this work.
Part of the work for this article has been conducted in the Research Campus MODAL funded by the German Federal Ministry of Education and Research (BMBF) (fund numbers 05M14ZAM, 05M20ZBM). GN is supported by ONR award # N000142312585. YC would like to acknowledge the support by the Quantum Engineering Program project A-8000339-01-00, by the Ministry of Education Singapore under project A-8000014-00-00, and by the National Research Foundation Singapore under its Industry Alignment Fund – Pre-positioning (IAF-PP) Funding Initiative and the Monetary Authority of Singapore under A-0003504-17-00. DVB would like to acknowledge the support by the Research Foundation Flanders (FWO) [23AXE35N].
@@
Appendix A Market Split Instance Generation
@@
    Generate a Random Solution Vector: Create a random solution vector $x$ where approximately half (±$(\pm 2)$ of the entries are set to $1$, and the rest are set to 00.
@@
      Identify a pair of entries in the row where one corresponds to 00 in $x$ and the other to $1$. Perform a “switch” that reduces the absolute slack of the row as much as possible.
@@
      Calculate the slack $s$ for the row. Set $c=\frac{3\cdot\text{numones}}{2}$, where num_ones is the number of $1$’s in the row. Find $c$ pairs of 00 and $1$, and adjust the corresponding entries by adding $\frac{s}{c}$ to the $1$’s and subtracting it from the 00’s.
@@
Appendix B Minimum Birkhoff Decomposition
@@
Next, we can map the Lehmer code to an integer in $[0,4!-1]$. In our example, the mapping of $[0,1,0,3]$ to an integer is as follows:
@@
Mapping an integer to $k$ unique integers. Suppose we want to map 4844 to $k=4$ unique integers. We start by finding the largest integer $x$ such that $\begin{pmatrix}x\\ 4\end{pmatrix}\leq 4844$. In this case, $x$ is equal to 19. Repeat the process for $9684844-\begin{pmatrix}19\\ 4\end{pmatrix}=$ 968 and $k-1$, i.e., find $x$ such that $\begin{pmatrix}x\\ 3\end{pmatrix}\leq 968$, which is 18. Repeating the process until the remainder is zero gives the integers $[19,18,17,16]$.
@@
Appendix C Maximum Independent Set Classical Post-processing
@@
$v\leftarrow\text{compute\_violations}(x)$triangleright map vertex to the number of constraints it violates
$c\leftarrow V$triangleright $c$ are vertices we try to add to the stable set
while sum($v$.values()) $>$do $\triangleright$ Make $x$ feasible
@@
while $\text{len}(c)>0$ $\triangleright$ Add vertices to the stable set and remain feasible

### v2

@@ 6 Conclusion and Discussion
It is widely believed that quantum computers can achieve an advantage over classical computers for some optimization problem instances in the future. While complexity theory suggests that, for NP-hard problems, the speed-up for finding provably optimal solutions is in general at most quadratic [13, 70, 63, 104], better quantum approximation algorithms with an exponential speed-up [129] or quantum optimization heuristics that outperform classical algorithms on certain instances may still exist. However, practical demonstrations of quantum advantage in optimization are still missing.
To identify problems where quantum advantage might be possible and to track progress toward achieving it, systematic analysis and benchmarking are key. To this end, we introduce [202], an open-source quantum optimization benchmarking library. We present ten classes of combinatorial optimization problems with instances of varying size and complexity for benchmarking classical and quantum optimization algorithms. Further, we define a reporting standard that ensures reproducibility and fair comparison while remaining flexible for evaluating different algorithms and hardware platforms.
Most problems are challenging for state-of-the-art classical solvers even at moderate size, and many of them are related to use cases of practical relevance. The problem classes exhibit diverse properties, such as different densities and types of constraints, to enable the study of different mathematical formulations and algorithmic approaches. This is essential for identifying problem structures suitable for quantum advantage in optimization.
We provide classical baseline results for all problem classes and quantum baseline results for selected ones to kick off this effort. However, performing exhaustive benchmarks is a massive endeavor and can only be achieved by the broader scientific community. Thus, we invite all researchers interested in advancing (quantum) optimization to investigate the provided problem instances and submit solutions using algorithms and hardware of their choice. We also want to emphasize the importance of improving classical methodologies, amongst other reasons, because quantum advantage can only be claimed if the best-known classical methods are outperformed. To accelerate the overall research, we also encourage researchers to report results for approaches that did not work well. To conclude, [202] is intended to be a living repository: novel solutions will be included on a rolling basis, and novel problem instances or adaptations of existing ones may be added as the state-of-the-art evolves.
@@ Acknowledgments
We want to thank Karen Aardal at TU-Delft for providing results for the Market Split instances using lattice point enumeration and Nils-Christian Kempke at ZIB for providing results by GPU-accelerated Schroeppel Shamir enumeration. We would like to thank Sébastien Designolle at the Zuse Institute Berlin for useful discussions on the Blended FW algorithm used in the classical benchmarking of the minimum Birkhoff decomposition problem (Section 4.3.4). We also thank Julien Gacon for valuable discussions and a helpful review, and Takashi Imamichi for assistance with software issues, sharing his insights, and providing feedback. Additionally, we are grateful to Andrew Wack and Paul Nation for insightful conversations about the benchmarking of quantum algorithms and their feedback on this work.
Part of the work for this article has been conducted in the Research Campus MODAL funded by the German Federal Ministry of Education and Research (BMBF) (fund numbers 05M14ZAM, 05M20ZBM, 05M2025). GN is supported by ONR award # N000142312585. YC would like to acknowledge the support from the Quantum Engineering Program project A-8000339-01-00, from the Ministry of Education Singapore under project A-8000014-00-00, and by the National Research Foundation Singapore under its Industry Alignment Fund – Pre-positioning (IAF-PP) Funding Initiative and the Monetary Authority of Singapore under A-0003504-17-00. DVB would like to acknowledge the support of the Research Foundation Flanders (FWO) [23AXE35N].
@@
Appendix A Building Capabilities via Instance Simplification
The problem instances we propose are challenging to even implement on current quantum computers due to noise, which limits the circuit depth. To track the progress of quantum algorithms and hardware today we therefore need a spectrum of simplified problem instances to observe a quantum signal. These instances will often be easy to solve classically. We propose to generate simplified QUBOs such that the time-evolution of the corresponding Hamiltonian becomes less difficult. More specifically, we consider a line of $n$ qubits which achieves all-to-all connectivity with a SWAP network with n2n-2 layers of SWAP gates [255]. Crucially, each entry $(i,j)$ of the QUBO matrix can be implemented after $d(i,j)\in\{0,...,n-2\}$ layers of SWAP gates. The distances $d_{i,j}$ allow us to define an $n\times n$ mask where entry $m_{i,j}(k)$ is one if $d(i,j)\leq k$ and zero otherwise. The number of allowed layers $k$ thus controls the difficulty of implementing $\exp(-iH_{C})$ on hardware. We simplify a QUBO matrix with this mask through the mapping $Q_{i,j}\to Q_{i,j}^{\prime}=m_{i,j}(k)Q_{i,j}$. Crucially, this only makes sense if the resulting QUBO $Q^{\prime}$ remains connected. In the benchmarking repository, we provide a code to simplify a QUBO matrix following this procedure. Crucially, in this process, the code generates a named field which contains the number of SWAP layers that constrain the QUBO for traceability.
Appendix B Market Split Classical Results
Fig 21 shows the results for Gurobi on a 32-core AMD EPYC-7542 processor using 8 cores and ABS2 [191] on an A100 SXM4 80GB, respectively, for different sizes of instances with different parameters $m$. ABS2 is very fast for smaller instances, but fails to find solutions within one hour for instances larger than size $m=6$.
Figure 21: Distribution of solution times using (top) Gurobi 11.0.0 with a time limit of 3600s and (bottom) ABS2 with a time limit of 600s for instances with different sizes $m$ and coefficients range $D$. All trials for $m\geq 8$ timed out for Gurobi and for $m\geq 7$ timed out for ABS2.
Appendix C Market Split Instance Generation
@@
    Generate a Random Solution Vector: Create a random solution vector $x$ where approximately half $(\pm 2)$ of the entries are set to $1$, and the rest are set to 0.
@@
      Identify a pair of entries in the row where one corresponds to 0 in $x$ and the other to $1$. Perform a “switch” that reduces the absolute slack of the row as much as possible.
@@
      Calculate the slack $s$ for the row. Set $c=\frac{3\cdot\text{numones}}{2}$, where num_ones is the number of $1$’s in the row. Find $c$ pairs of 0 and $1$, and adjust the corresponding entries by adding $\frac{s}{c}$ to the $1$’s and subtracting it from the 0’s.
@@
Appendix D Minimum Birkhoff Decomposition
@@
Next, we can map the Lehmer code to an integer in [0,4!1][0,4!-1]. In our example, the mapping of $[0,1,0,3]$ to an integer is as follows:
@@
Mapping an integer to $k$ unique integers. Suppose we want to map 4844 to $k=4$ unique integers. We start by finding the largest integer $x$ such that $\begin{pmatrix}x\\ 4\end{pmatrix}\leq 4844$. In this case, $x$ is equal to 19. Repeat the process for $4844-\begin{pmatrix}19\\ 4\end{pmatrix}=968$ and k1k-1, i.e., find $x$ such that $\begin{pmatrix}x\\ 3\end{pmatrix}\leq 968$, which is 18. Repeating the process until the remainder is zero gives the integers $[19,18,17,16]$.
@@
Appendix E Maximum Independent Set Classical Post-processing
@@
$v\leftarrow\text{compute\_violations}(x)$ $\triangleright$ map vertex to the number of constraints it violates
$\leftarrow V \triangleright cc$ are vertices we try to add to the stable set
while sum($v$.values()) >>do $\triangleright$ Make $x$ feasible
@@
[…]